# Sculpting Molecules in 3D: A Flexible Substructure Aware Framework for Text-Oriented Molecular Optimization

### v1

@@ 2.5 Hard-coded molecule optimization on appointed sites
From a methodological perspective, the 3D positions of the appointed sites are utilized during the hard-coded optimization process (see Supplementary Section S4 for the algorithm details). We showcase 3DToMolo’s capability on two exemplary drug-related molecules, penicillin and triptolide. For penicillin, the crucial $\beta$-lactam ring [49, 50] is vulnerable to $\beta$-lactamase binding [51] and acidic hydrolysis [52] (Figure 6(a)). One proposed mechanism suggests the initial nucleophilic attack by the oxygen atom from another amide group on the carbonyl group in the $\beta$-lactam ring [51]. Thus, an effective strategy is replacing the benzyl group by a more electron-withdrawing functional group with substantial steric volume to impede lactamase binding and to weaken the nucleophilicity of the attacking oxygen atom. We optimize the penicillin molecule under the prompt "This molecule has large electron-withdrawing groups" while maintaining structural constraints on the entire molecule except for the benzyl group. Of the optimized structures, 23% successfully exhibit a decrease in the electron density on the attacking oxygen atom revealed by DFT computation, with 43% of these structures featuring a ring of at least five members, indicative of substantial steric effects. An exemplary result demonstrates the replacement of the benzyl group with an isoxazole group (Figures 6(b) and 6(c)). Notably, the isoxazolyl series of semisynthetic penicillins, such as oxacillin, has been recognized for superior resistance to acids and $\beta$-lactamases [53]. A comparative test utilizing GPT-3.5 as a baseline reveals its inability to generate valid SMILES strings or preserve the core structure under varying instructions (Supplementary Section S9).
@@
Figure 6: Case studies for molecule optimization on appointed sites. (a) The molecular structure of penicillin. The blue line marks the $\beta$-lactam ring. The dashed arrow shows the nucleophilic attack from the oxygen atom on the side chain to the $\beta$-lactam ring. The yellow region marks the functional group to be optimized. (b) One representative example of our optimization results under the given prompt is shown in the figure. The green region marks the modification where the benzyl group is replaced by a more electron withdrawing isoxazole group. (c) The three-dimensional structures of the penicillin molecule (top) and the isoxazole-replacing derivative (below) energy-optimized by Density Functional Theory (DFT) computation. The numbers are the partial charge on the attacking oxygen atom, showing that our optimization has successfully reduced its electron density and its nucleophilicity. (d) The molecular structure of triptolide. The yellow regions mark the optimization sites. (e) Two exemplary optimizations where the two modified side chains have connected. Both exhibit increases in solubility as well as in polarity, as the given prompt orders.
@@
We further demonstrate the capability of 3DToMolo beyond simple side chain substitution in the optimization of Triptolide [54, 55] with the goal of enhancing its water solubility. Referencing to reported modified derivatives [56], we choose two adjacent sites for optimization, as depicted in Figure 6(d). The remaining structure is constrained. We employ the prompt "This molecule is soluble in water and has high polarity" that has been proven to be more effective in the previous section. Within successful optimizations, we observe some connected structures between these two optimization sites (Figure 6(e)), which is not achievable through sequential side chain substitution. These findings underscore the remarkable proficiency of 3DToMolo in selectively appointing and modifying substructures based on natural language guidance, particularly in scenarios involving complex isomeric structures and three-dimensional considerations.
@@ 3 Discussion
Considering the crucial aspect of data efficiency, our design involves the decoupling of the molecule structure generation model and the text-structural alignment guidance. This separation enables us to leverage vast amounts of unlabeled structure data during the training of the structure generator. 3DToMolo proves particularly advantageous when labeled data for guiding text-structural alignment is limited. By tapping into the abundance of unlabeled data, our model gains a robust understanding of diverse molecular structures. Remarkably, we are able to selectively fine-tune specific aspects of the model in a low-rank manner [57], particularly when dealing with intricate molecular geometry configurations such as binding scenarios or conformers exhibiting high energy states. An additional rationale for favoring diffusion models in molecule optimization lies in their global optimization approach. This stands in contrast to the local reinforced, and often greedy approach of optimizing one disconnection site at a time. The global methodology employed by diffusion models contributes to the generation of more diverse and comprehensive optimization results. Finally, it is noteworthy that 3DToMolo distinguishes itself by not necessitating optimization trajectories as part of our training data, in contrast to traditional MCTS methods. This is attributed to the fixed nature of the noising process equation for every molecule, and only the denoising process is learned. Conversely, the manipulation of the noising process to reinforce intermediate states for synthesizability (as an example) is a plausible avenue. This adjustment holds the potential to be beneficial for generating retrosynthetic pathways for our optimized molecules.
@@
As the field of multi-modality large neural networks advances swiftly, our research is only a preliminary attempt on harnessing the potential of multimodality information to guide the optimization of molecular structures. As an illustrative example, we utilized paired text-molecule data to train the alignment between the molecular representation $X$ and the corresponding text representation $Y$. What remains unexplored is the potential of adversarial matching between $X$ and $Y$, wherein the mapping function $G$ learns to map the distribution of $X$ to that of $Y$. This approach has the capacity to leverage unpaired text and molecule data, offering a promising avenue for further investigation. On the text side, our approach involves utilizing a pretrained Large Language Model (LLM) for extracting text embeddings. This LLM model is specifically trained by predicting the next word token based on context. While effective, a more intricate strategy involves directly training a text-molecule equivariant large model in a similar way as Emu2 [58]. This advanced approach allows for the generation of multi-modality outputs. Unlike our current approach, where the model is trained to predict molecules given the text, this more involved method also operates in reverse. It trains the model not only to predict molecules based on textual descriptions but also to predict text from molecular information. This bidirectional training scheme contributes to a more versatile and expressive representation. Moreover, extending beyond text, the inclusion of illustrations from academic papers adds another layer of informative guidance for optimizing molecules. Such image representations can also serve as valuable cues for refining and enhancing the structural optimization process.
@@ 4.1 Datasets
We use PCQM4Mv2 dataset  [^1] https://ogb.stanford.edu/docs/lsc/pcqm4mv2/ to pretrain an unconditional diffusion model for modeling complex data distribution to generate new structures within the chosen chemical space. PCQM4Mv2 is a quantum chemistry dataset including 3,746,619 molecules originating from the PubChemQC project [61]. MoleculeSTM dataset [11] with over 280K chemical structure-text pairs is used to train a text-molecule model. To better align chemical space with semantic space, we effectively incorporate 3D structure information of molecules to enhance the alignment with textual description. However, MoleculeSTM lacks 3D coordinates of molecules, thus we extract 3D information and energy-related values from PubchemQC according to the PIDs in MoleculeSTM. In terms of downstream tasks, a novel molecule could be generated from Gaussian noise, or a molecule selected from Zinc dataset [42] could be optimized through applying noise and recursively denoising.
@@ 4.3.1 Denoising diffusion process
In the first stage, we conduct pretraining of a generative 2D-3D molecular diffusion model. Following the structure of typical diffusion models, our model encompasses two key processes: the forward process and the reverse process. To adapt these processes to 2D-3D molecular graphs, we represent the molecule $M$ as a combination of node features (atom types) $H$, an adjacency matrix $E$ (representing chemical bond edges), and 3D positions $P$, denoted as $(H,E,P)$. Specifically, suppose the molecule is composed of $n$ atoms, then $H=[H^{1},\dots,H^{n}]$ represents the one-hot embedding in the periodic table for the $n$ atoms. In the forward process, the original structure of the molecule undergoes a joint Markovian transition step by step. We denote the intermediate structure as $M_{t}:=(H_{t},E_{t},P_{t})$, where the initial structure is denoted as $M_{0}$. For the 3D point cloud $P_{t}$, a Markov chain is implemented by incrementally introducing Gaussian noise over $T$ steps. The transition from $P_{t-1}$ to $P_{t}$ is described as follows:

### v2

@@ 2.5 Hard-coded molecule optimization on appointed sites
From a methodological perspective, the 3D positions of the appointed sites are utilized during the hard-coded optimization process (see Supplementary Section S4 for the algorithm details). We showcase 3DToMolo’s capability on two exemplary drug-related molecules, penicillin and triptolide. For penicillin, the crucial $\beta$-lactam ring [55, 56] is vulnerable to $\beta$-lactamase binding [57] and acidic hydrolysis [58] (Figure 6(a)). One proposed mechanism suggests the initial nucleophilic attack by the oxygen atom from another amide group on the carbonyl group in the $\beta$-lactam ring [57]. Thus, an effective strategy is replacing the benzyl group by a more electron-withdrawing functional group with substantial steric volume to impede lactamase binding and to weaken the nucleophilicity of the attacking oxygen atom. We optimize the penicillin molecule under the prompt "This molecule has large electron-withdrawing groups" while maintaining structural constraints on the entire molecule except for the benzyl group. Of the optimized structures, 23% successfully exhibit a decrease in the electron density on the attacking oxygen atom revealed by DFT computation, with 43% of these structures featuring a ring of at least five members, indicative of substantial steric effects. An exemplary result demonstrates the replacement of the benzyl group with an isoxazole group (Figures 6(b)). Notably, the isoxazolyl series of semi-synthetic penicillins, such as oxacillin and (Figure 6(c)), has been recognized for superior resistance to acids and $\beta$-lactamases [59]. A comparative test utilizing GPT-3.5 as a baseline reveals its inability to generate valid SMILES strings or preserve the core structure under varying instructions (Supplementary Section S10).
@@
Figure 6: Drug molecule optimization on appointed sites. (a) The molecular structure of penicillin. The blue line marks the $\beta$-lactam ring. The dashed arrow shows the nucleophilic attack from the oxygen atom on the side chain to the $\beta$-lactam ring. The yellow region marks the functional group to be optimized. The partial charge on the nucleophilic-attacking oxygen atom is -0.506. (b) One representative example of our optimization results under the given prompt is shown in the figure. The green region marks the modification where the benzyl group is replaced by a more electron withdrawing isoxazole group. With the modification, the partial charge on the nucleophilic-attacking oxygen atom is successfully reduced to -0.488. (c) The molecular structure of oxacillin, an existing analogue to our modification. (d) The molecular structure of triptolide. The yellow regions mark the optimization sites. (e) Two exemplary optimizations where the two modified side chains have connected. Both exhibit increases in solubility as well as in polarity, as the given prompt orders. (f) Selected examples in diverse drug molecule optimization tasks that have reported analogs.
@@
We further demonstrate the capability of 3DToMolo beyond simple side chain substitution in the optimization of Triptolide [60, 61] with the goal of enhancing its water solubility. Referencing to reported modified derivatives [62], we choose two adjacent sites for optimization, as depicted in Figure 6(d). The remaining structure is constrained. We employ the prompt "This molecule is soluble in water and has high polarity" that has been proven to be more effective in the previous section. Within successful optimizations, we observe some connected structures between these two optimization sites (Figure 6(e)), which is not achievable through sequential side chain substitution.
To provide a quantitative comparison, we construct a small task set of optimization on appointed sites in common drug molecules under given prompts. Akin to penicillin and triptolide above, modification sites and optimization strategies of each molecule are carefully chosen based on relevant studies (details in Supplementary Section S2 and Figure S1). The performance of 3DToMolo is summarized in Table 4, whereas GPT-3.5 is used as a baseline. Compared to the failure of GPT-3.5 shown in the table, 3DToMolo is more powerful in adapting the domain knowledge it has learned in diverse drug discovery tasks. Several optimized results by 3DToMolo not only meet the structural and property requirements, but also are chemically reasonable, as they resemble known molecules in existing database (Figure 6(f)). These findings underscore the remarkable proficiency of 3DToMolo in selectively appointing and modifying substructures based on natural language guidance, particularly in scenarios involving complex isomeric structures and three-dimensional considerations.
Table 4: The results of molecule optimization based on specified editing positions on the small molecule drug dataset, with the evaluation measured by the hit ratio (%) of the property changes in each objective-oriented prompt. Statistically significant improvement (t-test over 5 different dataset splits, p-value$<$ 0.05) is highlighted with bold text.
Drug Molecule Objectives Prompts Validity (%) Hit Ratio (%)
GPT-3.5 3DToMolo GPT-3.5 3DToMolo
Apixaban Electron-donating substitution High HOMO energy 55 73 1 18
Empagliflozin More hydrophilic Large polarity 0 15 0 8
Ibrutinib Higher permeability High permeability 19 33 0 2
Dapagliflozin Large aqueous solubility More soluble 0 35 0 13
Osimertinib Electron-withdrawing substitution Low HOMO energy 30 52 0 39
Olaparib Electron-withdrawing substitution Low HOMO energy 34 30 1 2
Abemaciclib More hydrogen bonding interactions More hydrogen bond acceptors 20 42 0 21
More hydrogen bond donors 0 45 0 25
Pomalidomide Electron-withdrawing substitution Low HOMO energy 24 64 11 10
Tafamidis More hydrogen bonding interactions More hydrogen bond acceptors 46 61 0 13
More hydrogen bond donors 10 48 0 5
Lisdexamfetamine To control the hydrolysis rate High HOMO energy 0 11 0 8
Low HOMO energy 0 19 0 5
  • Note that the prompts shown in the table have been abbreviated. The prompts actually used in 3DToMolo can be found in Table 2, while we formulated our request to GPT-3.5 as follows: "Optimize the following molecule: <SMILES> to achieve the desired properties: <specified properties>, while ensuring that the specified substructure <substructure> remains unchanged. Provide 10 different and valid results." Note in this experimental setting, we employed a multi-run approach to facilitate the generation of a total of 100 optimized results for each drug molecule. Specifically, each iteration obtained 10 distinct outputs from GPT-3.5, with a total of 10 iterations conducted. In most cases, GPT-3.5 fails to meet the requirements, as the generated molecules do not conform to either the specified structural criteria or the desired properties.
@@ 3 Discussion
Considering the crucial aspect of data efficiency, our design involves the decoupling of the molecule structure generation model and the text-structural alignment guidance. This separation enables us to leverage vast amounts of unlabeled structure data during the training of the structure generator. 3DToMolo proves particularly advantageous when labeled data for guiding text-structural alignment is limited. By tapping into the abundance of unlabeled data, 3DToMolo gains a robust understanding of diverse molecular structures. Remarkably, we are able to selectively fine-tune specific aspects of the model in a low-rank manner [63], particularly when dealing with intricate molecular geometry configurations such as binding scenarios or conformers exhibiting high energy states. An additional rationale for favoring diffusion models in molecule optimization lies in their global optimization approach. This stands in contrast to the local reinforced, and often greedy approach of optimizing one disconnection site at a time. The global methodology employed by diffusion models contributes to the generation of more diverse and comprehensive optimization results. Finally, it is noteworthy that 3DToMolo distinguishes itself by not necessitating optimization trajectories as part of our training data, in contrast to traditional MCTS methods. This is attributed to the fixed nature of the noising process equation for every molecule, and only the denoising process is learned. Conversely, the manipulation of the noising process to reinforce intermediate states for synthesizability (as an example) is a plausible avenue. This adjustment holds the potential to be beneficial for generating retrosynthetic pathways for our optimized molecules.
@@
As the field of multi-modality large neural networks advances swiftly, our research is only a preliminary attempt on harnessing the potential of multimodality information to guide the optimization of molecular structures. As an illustrative example, we utilized paired text-molecule data to train the alignment between the molecular representation $X$ and the corresponding text representation $Y$. What remains unexplored is the potential of adversarial matching between $X$ and $Y$, wherein the mapping function $G$ learns to map the distribution of $X$ to that of $Y$. This approach has the capacity to leverage unpaired text and molecule data, offering a promising avenue for further investigation. On the text side, 3DToMolo involves utilizing a pretrained Large Language Model (LLM) for extracting text embeddings. This LLM model is specifically trained by predicting the next word token based on context. While effective, a more intricate strategy involves directly training a text-molecule equivariant large model in a similar way as Emu2 [64]. This advanced approach allows for the generation of multi-modality outputs. Unlike our current approach, where the model is trained to predict molecules given the text, this more involved method also operates in reverse. It trains the model not only to predict molecules based on textual descriptions but also to predict text from molecular information. This bidirectional training scheme contributes to a more versatile and expressive representation. Moreover, extending beyond text, the inclusion of illustrations from academic papers adds another layer of informative guidance for optimizing molecules. Such image representations can also serve as valuable cues for refining and enhancing the structural optimization process.
@@ 4.1 Datasets
We use PCQM4Mv2 dataset  [^1] https://ogb.stanford.edu/docs/lsc/pcqm4mv2/ to pretrain an unconditional diffusion model for modeling complex data distribution to generate new structures within the chosen chemical space. PCQM4Mv2 is a quantum chemistry dataset including 3,746,619 molecules originating from the PubChemQC project [67]. MoleculeSTM dataset [11] with over 280K chemical structure-text pairs is used to train a text-molecule model. To better align chemical space with semantic space, we effectively incorporate 3D structure information of molecules to enhance the alignment with textual description. However, MoleculeSTM lacks 3D coordinates of molecules, thus we extract 3D information and energy-related values from PubchemQC according to the PIDs in MoleculeSTM. Regarding downstream tasks, a novel molecule could be generated from Gaussian noise, or a molecule selected from Zinc dataset [47] could be optimized through applying noise and recursively denoising.
@@ 4.3.1 Denoising diffusion process
In the first stage, we conduct pretraining of a generative 2D-3D molecular diffusion model. Following the structure of typical diffusion models, 3DToMolo encompasses two key processes: the forward process and the reverse process. To adapt these processes to 2D-3D molecular graphs, we represent the molecule $M$ as a combination of node features (atom types) $H$, an adjacency matrix $E$ (representing chemical bond edges), and 3D positions $P$, denoted as $(H,E,P)$. Specifically, suppose the molecule is composed of $n$ atoms, then $H=[H^{1},\dots,H^{n}]$ represents the one-hot embedding in the periodic table for the $n$ atoms. In the forward process, the original structure of the molecule undergoes a joint Markovian transition step by step. We denote the intermediate structure as $M_{t}:=(H_{t},E_{t},P_{t})$, where the initial structure is denoted as $M_{0}$. For the 3D point cloud $P_{t}$, a Markov chain is implemented by incrementally introducing Gaussian noise over $T$ steps. The transition from $P_{t-1}$ to $P_{t}$ is described as follows:
@@ 4.3.2 Text-Structure Alignment
It is worth emphasizing that all the aforementioned optimization experiments were conducted with a focus on adapting the de-noising steps. Our findings from extra experiments (detailed in Supplementary Section S6) and inherent nature of diffusion model indicate that the choice of de-noising steps is correlated with the resemblance between the optimized molecule and the initial input. In cases where minimal alterations to the input were anticipated, opting for smaller de-noising steps proved to be effective.
@@ 4.4 Evaluation of physiochemical properties
Author Contributions
K. Z. and W. D. conceptualized the research. K. Z. wrote the code. K. Z., W. D., and Y. L. designed the downstream tasks. K. Z., Y. L., and G. W. analyzed the data and results. K. Z., W. D., and Y. L. drafted the manuscript. B. W., Y. R., and X.-Y. Z. supervised the experimental process.
Competing Interests
The authors declare no competing interest.
Data Availability
All datasets used in this document are publicly available. The PCQM4MV2 dataset including optimized structures, various properties and 3D information of all the 3,378,606 training molecules is available at acs.jcim.7b00083: https://doi.org/10.1021/acs.jcim.7b00083. The MoleculeNet dataset is available at C7SC02664A: https://doi.org/10.1039/C7SC02664A. The MoleculeSTM dataset is available at s42256-023-00759-6: https://doi.org/10.1038/s42256-023-00759-6. The PubChemQC dataset is available from https://nakatamaho.riken.jp/pubchemqc.riken.jp/. We used the HOMO, LUMO, and HOMO-LUMO gap from PubChemQC dataset, which was obtained at B3LYP/6-31G* level. The redox dataset is available in the Supplementary of acsomega.8b00576: https://doi.org/10.1021/acsomega.8b00576.
5 Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) (Grant 62376265). And we thank Chunngai Hui for the insightful discussions in organic chemistry and small-molecule drug related tasks.